Qd = 180 – 2P


Qs = - 15 + P


If a tax of Rs 2/unit is imposed on sellers, calculate the effect on the price paid by buyers


and price received by sellers. Demonstrate that the effect would be unchanged if the tax


had been imposed on buyers instead of sellers. (2+2.75 +3)

1
Expert's answer
2022-02-28T09:39:23-0500

"Q_d = 180 \u2013 2P"

"P = \frac{180 – Q_d}{2}"

"Q_s = - 15 + P"

"P= Q_s+15"

Pretax equilibrium

"180 \u2013 2P=- 15 + P"

"3P= 195"

P= 65

Q*= -15+ 65= 50

After Tax

"P= Q_s+17"

"\therefore Q+17= \frac{180 – Q}{2}"

"2Q+34=180-Q"

3Q= 146

Q= 49

Suppliers receive in pretax supply

P= 15+ 49

= 64

Consumers receive the following

"P = \frac{180 – 49}{2}= \frac{131}{2}= 65.5"

If tax is imposed on the consumers

"P_T= 90-0.5Q+2= 98-0.5Q"

"98-0.5Q= 15+Q"

1.5Q= 83

Q= 53.33

Price for consumers pretax supply

90- 0.5(53.33)= 63.5

Price for suppliers

P= 15+53= 78

NB: The conclusion is that whether the tax is imposed on the consumer or suppliers the effect is the same.